A boy throws aball in air in such a manner that when the ball is in its max height, he throws a anothervball after a time diif of 1 sec. What will be the height atained by them

Grade:12th pass

1 Answers

Vikas TU
7 years ago
Time difference that is 1 sec.
That means the time of flight is 2 sec.
Hence,
T = 2 = 2u/g
u = 10 m/s.
H =u^2/2g = > 100/20 = 5 meter.